OpenStudy (anonymous):

Simplify 4 square root of 98 28 square root of 2 7 square root of 2 196 square root of 2 11 square root of 2

OpenStudy (anonymous):

@ihateschool18 @zaibali.qasmi

OpenStudy (here_to_help15):

:P

OpenStudy (here_to_help15):

is that the full question?

OpenStudy (anonymous):

yess

OpenStudy (here_to_help15):

im trying to help 4 ppl at once hold on just 1 sec :P

OpenStudy (anonymous):

I believe It would be A

OpenStudy (anonymous):

I believe It would be A\[\sqrt{98}=7\sqrt{2}\] and 7x4=28 so \[28\]

OpenStudy (anonymous):

\[28\sqrt{2}\]***